# Neuro-Symbolic Control with Large Language Models for Language-Guided Spatial Tasks

### v1

@@ 4.1.2 Task Specification
We evaluate four canonical spatial relations constituting the task space $\Omega$:
$\Omega=\{\texttt{right\_of},\texttt{left\_of},\texttt{above},\texttt{below}\}.$ (35)
A task $\mathcal{T}\in\Omega$ is considered successfully completed if the corresponding geometric constraint is satisfied with margin $m=50$ pixels:
$\mathcal{G}(\mathbf{s}_{t},\mathcal{T})=1,\quad\text{for some }t\leq T.$ (36)
Each task is presented to the system via natural language instruction, requiring interpretation by the symbolic reasoning module.
@@ 4.1.3 Compared Methods
To isolate the contributions of symbolic reasoning and neural execution, we evaluate three control paradigms:
@@
In this baseline, a large language model directly predicts absolute target coordinates $(\hat{x}_{r},\hat{y}_{r})$ at each time step based on the current state $\mathbf{s}_{t}$ and natural language instruction $\mathcal{T}_{\text{nl}}$. The displacement is computed as:
$(\Delta x_{t},\Delta y_{t})=(\hat{x}_{r}-x_{r}^{t},\hat{y}_{r}-y_{r}^{t}).$ (37)
No learned motion prior is applied. This baseline evaluates the capacity of LLMs to perform spatial reasoning and continuous control without neural assistance.
@@
This baseline employs the neural delta controller exclusively with a fixed, ground-truth task encoding:
$(\Delta x_{t},\Delta y_{t})=f_{\theta}(\mathbf{u}_{t}),\quad\phi(\mathcal{T})=\phi^{\ast},$ (38)
where $\phi^{\ast}$ denotes the oracle task encoding. No LLM reasoning is invoked during execution. This configuration evaluates the controller’s performance when task semantics are perfectly specified, establishing an upper bound on neural execution quality.
@@
The proposed neuro-symbolic method combines symbolic reasoning with neural execution:
$(\Delta x_{t},\Delta y_{t})=f_{\theta}\left(\mathbf{u}_{t},\pi_{\text{sym}}(\mathbf{s}_{t},\mathcal{T})\right).$ (39)
A local LLM interprets the natural language instruction and outputs a symbolic relation label $\mathbf{z}_{t}\in\mathcal{Z}$, which conditions the neural delta controller. The controller predicts incremental displacements in a closed-loop manner, enabling stable and efficient convergence.
4.1.4 Language Models
All symbolic reasoning is performed using locally deployed large language models to ensure reproducibility, low-latency inference, and independence from external APIs. We evaluate the following models:
  • Mistral-7B: A 7-billion parameter model optimized for instruction following and efficient inference.
  • Phi-2: A compact 2.7-billion parameter model designed for resource-constrained deployment.
  • LLaMA-3.2: Meta’s open-weight model with strong reasoning capabilities.
The neural controller $f_{\theta}$ remains fixed across all experiments, enabling controlled comparison of symbolic reasoning quality across models while isolating the contribution of the neural execution layer.
4.1.5 Evaluation Protocol
For each combination of this proposed method, language model, and task, we conduct $N=20$ independent episodes with randomized initial configurations. We showed mean performance with standard deviation to quantify variability. In Fig. 3, we can see that the proposed LLM+DL framework consistently achieves higher success rates across all spatial tasks when compared to both LLM-only and DL-only baselines.
@@ 4.1.6 Evaluation Metrics
We assess performance using the following metrics:
@@ 4.2 Results
4.2.1 Overall Performance Comparison
@@
Several observations emerge from the results:
  • For the consistent improvement, across all language models, the suggested LLM+DL framework produces success rates that are either greater or equivalent to those of LLM-only control, with significant drops in average steps.
  • For the efficiency gains, depending on the underlying language model, the neural delta controller allows convergence in many fewer steps, with speedups ranging from $2.52\times$ to $5.82\times$.
  • For the compensation for weak reasoning, the Phi model shows the best LLM+DL success rate (0.85) but the lowest LLM-only success rate (0.60), indicating that the neural controller successfully makes up for weaker symbolic reasoning.
4.2.2 Convergence Analysis
According to Fig. 4, LLM+DL typically reduces the number of necessary control steps by more than 70%.
@@ 4.2.2 Convergence Analysis
Each method’s unique qualities are shown by the convergence behavior:
@@
The normalized distance-to-goal over time is shown in Fig. 5, which highlights the LLM+DL framework’s faster and more stable convergence.
Figure 5: Normalized distance-to-goal over time for the right_of task. LLM+DL exhibits fast, monotonic convergence with low variance, whereas LLM-only control shows slower and less stable behavior.
4.2.3 Relative Improvement Analysis
@@
Key findings include:
  • For the significant speedup, for all tasks, the suggested method delivers an average speedup of $5.58\times$, with a maximum speedup of $8.83\times$ for the left_of relation.
  • For the task-dependent gains, compared to orthogonal relations (right_of, above), lateral relations (left_of, below) show greater increases in speedup and success rate ($\Delta\text{SR}=0.48$). This imbalance implies that some spatial interactions are more difficult for LLM-only management because of incorrect or hallucinated coordinate predictions.
  • Regardless of the kind of work, step reduction consistently surpasses 70%, indicating that the neural delta controller significantly increases control efficiency.
As summarized in Fig. 6, the proposed method achieves speedups of up to $8.83\times$ over LLM-only control.
@@
4.2.4 Robustness Across Language Models
Performance gains are consistent across all evaluated language models, with several notable observations:
@@ 4.2.4 Robustness Across Language Models
  • To the decreased variance, the LLM+DL performs more consistently across episodes and beginning settings, as seen by a significantly lower standard deviation of success rates than LLM-only.
@@ 4.2.5 Ablation: Component Contribution
5 Discussion and Conclusion
The experimental findings show that the efficiency, stability, and resilience of language-conditioned control significantly increase when symbolic reasoning and continuous control are explicitly separated. The suggested neuro-symbolic architecture avoids the instability, sluggish convergence, and hallucinated actions frequently seen in monolithic LLM-based control systems by assigning high-level semantic interpretation to a big language model and low-level execution to a neural delta controller.
The suggested LLM+DL method continuously beats LLM-only baselines in both success rate and convergence speed across all assessed spatial tasks and local language models. While preserving or increasing task success rates, average step reductions surpass 70%, with speedups up to $8.83\times$ in some setups. Interestingly, these improvements hold true even with poorer or smaller language models, suggesting that the neural execution layer successfully makes up for flawed symbolic thinking.
In order to stabilize execution, the neural delta controller is essential. It demonstrates the efficacy of learning task-agnostic motion primitives by generalizing across tasks and language models without retraining after being trained exclusively on synthetic geometric data. Unlike opaque end-to-end LLM control, restricting the LLM to symbolic outputs allows failures to be clearly attributed to either semantic reasoning or execution, which further improves interpretability and debuggability.
Practically speaking, the framework provides a number of deployment benefits. Real-time execution is made possible by the lightweight neural controller, independent component updates are made possible by the modular architecture, and local LLM inference minimizes external dependencies. Because of these characteristics, the method works effectively in environments where resources are limited and safety is crucial.
The present analysis is restricted to planar environments with perfect state observability and primitive spatial linkages, despite these advantages. The expressiveness of compositional or temporally extended instructions is limited by the symbolic layer’s operation over a fixed task set. Furthermore, even if local LLMs lower reliance and cost, stochastic outputs and inference latency are still factors to take into account.
This framework will be expanded in the future to include real-world robotic platforms, richer symbolic representations, hierarchical and compositional task structures, and multi-object and vision-based environments. One interesting way to scale the technique is to incorporate long-horizon planning and perception-driven state estimation.
To sum up, our research gives compelling empirical proof that neuro-symbolic integration provides a rational and practical means of fusing language comprehension with ongoing control. The suggested framework promotes the creation of dependable, comprehensible, and effective language-guided embodied systems by utilizing the complementing capabilities of big language models and neural controllers.
@@ 5 Discussion and Conclusion
Figure 10: Normalized distance-to-goal over time for all spatial tasks. The proposed LLM+DL framework converges faster and more stably than LLM-only control across all relations.
@@
The efficacy of clearly distinguishing symbolic reasoning from continuous execution in language-conditioned control tasks is further supported by these results.

### v2

@@ 4.1.1 Task Specification
We evaluate four canonical spatial relations: $\{\texttt{right\_of},\texttt{left\_of},\texttt{above},\texttt{below}\}$. A task is considered successful if the corresponding spatial constraint is satisfied within a tolerance margin of $m=50$ pixels at any time step $t\leq T$.
Each task is provided to the system in natural language form. For example, the instruction “Move the red object to the right of the blue object” corresponds to the right_of relation. The same task set and margin values are used across all compared control strategies.
@@ 4.1.2 Compared Methods
We compare three control strategies to isolate the contribution of symbolic reasoning and neural execution.
@@
In this baseline, a large language model directly predicts absolute target coordinates at each control step based on the current state and the natural language instruction. The resulting displacement is computed as the difference between the predicted and current target positions. No learned motion prior or bounded control mechanism is applied.
@@
This baseline uses the neural delta controller exclusively, conditioned on a fixed oracle task encoding corresponding to the ground-truth spatial relation. No language model inference is performed during execution. This configuration serves as an upper bound on execution performance when task semantics are perfectly specified.
@@
The proposed method combines symbolic reasoning and neural execution. At each control step, a local language model infers a symbolic task label from the natural language instruction and current state. This symbolic output conditions the neural delta controller, which generates bounded incremental displacements in a closed-loop manner.
All methods share the same environment, control horizon, and evaluation protocol.
4.1.3 Large Language Models
Symbolic reasoning is performed using locally deployed large language models to ensure reproducibility, low-latency inference, and independence from external APIs. The following models are evaluated:
  • Mistral-7B, a 7B-parameter instruction-tuned model,
  • Phi-2, a 2.7B-parameter compact model optimized for efficiency,
  • LLaMA-3.2, an open-weight model with strong general reasoning capability.
The neural delta controller $f_{\theta}$ is kept fixed across all experiments. This design isolates the effect of symbolic reasoning quality and ensures that performance differences arise from language model inference rather than changes in low-level control dynamics.
@@ 4.1.4 Evaluation Metrics
For each combination of our proposed method, language model, and also the task, we conduct a comprehensive $N=20$ independent episode with the randomized initial configurations to show the strong results. Also, we showed the mean performance with standard deviation to quantify variability of the proposed framework.
We assess performance using the metrics like success rate [49], average steps and distance to goal [50]:
@@ 4.2 Results
In this section, we discussed the results obtained on different LLM models and what are the key improvements and observations emerged from these experiments. In Figure 3, we can see that the proposed LLM+DL framework consistently achieves higher success rates across all the spatial tasks when it is compared to both LLM-only and DL-only baselines frameworks.
@@
Several key observations are emerged from the results. Among which, the first one is, for the consistent improvement, across all language models, the suggested LLM+DL framework produces success rates that are either greater or equivalent to those of LLM-only control, with significant drops in average steps. Secondly, for the efficiency gains, depending on the underlying language model, the neural delta controller allows convergence in many fewer steps, with speedups ranging from $2.52\times$ to $5.82\times$. The third most significant observation was for the compensation for weak reasoning for which the Phi model shows the best LLM+DL success rate (0.85) but the lowest LLM-only success rate (0.60), indicating that the neural controller successfully makes up for weaker symbolic reasoning.
According to Figure 4, LLM+DL typically reduces the number of necessary control steps by more than 70%. Each method’s unique qualities are shown by the convergence behavior:
@@
The normalized distance-to-goal over time is shown in Figure 5, which highlights the LLM+DL framework’s faster and more stable convergence.
Figure 5: Normalized distance-to-goal over time for the right_of task. LLM+DL exhibits fast, monotonic convergence with low variance, whereas the LLM-only control shows slower and less stable behavior.
@@
For the significant speedup, for all tasks, the suggested method delivers an average speedup of $5.58\times$, with a maximum speedup of $8.83\times$ for the left_of relation. For the task-dependent gains, compared to orthogonal relations (right_of, above), lateral relations (left_of, below) show greater increases in speedup and success rate ($\Delta\text{SR}=0.48$). This imbalance implies that some spatial interactions are more difficult for LLM-only management because of incorrect or hallucinated coordinate predictions. Regardless of the kind of work, step reduction consistently surpasses 70%, indicating that the neural delta controller significantly increases control efficiency.
As summarized in Figure 6, the proposed method achieves speedups of up to $8.83\times$ over LLM-only control.
As the performance gains are consistent across with several notable observations:
@@
  • Due to the decreased variance, the LLM+DL performs more consistently across episodes and beginning settings, as seen by a significantly lower standard deviation of success rates than LLM-only.
@@ 4.2.1 Ablation: Component Contribution
5 Discussion
5.1 Deterministic Stability vs Motion Fluidity
The symbolic discretization mainly prioritizes deterministic and stable execution by preventing hallucinated actions, but it may produce less fluid trajectories in some cases. This trade-off is intentional for the purpose of safety and convergence, and the framework can further be extended to finer symbolic resolutions (e.g., headings or intensities) while preserving bounded control.
5.2 Semantic Limits of Local Language Models
Locally deployed language models may introduce a lower semantic ceiling than cloud-scale options, specifically for abstract or compositional instructions. Since the neural controller cannot correct incorrect symbolic intent, but it does stabilize execution once a symbolic decision is made, and richer symbolic reasoning can be incorporated as per the need.
5.3 Robustness to Sensor Noise and State Uncertainty
The evaluation presented in the paper assumes ground-truth state information to isolate control behavior. In the presence of moderate sensor noise, the closed-loop delta controller can modify and correct transient symbolic errors as state estimates are updated, while persistent uncertainty would require additional perceptual filtering.
5.4 Generalization Beyond 2D Environments
Although evaluated in a planar 2D setting, the neuro-symbolic decomposition presented in this paper is not inherently restricted to only two dimensions. Extension to 3D environments would involve increasing the state representation and symbolic predicate set. Irregular object shapes can be accommodated through geometry-aware symbolic abstractions. This modular design facilitates such extensions without changing the core execution controller.
6 Conclusion
In this paper a neuro-symbolic control framework for language-guided spatial manipulation that explicitly separates high-level symbolic reasoning from low-level continuous execution is presented. By constraining the language model to symbolic task outputs and delegating numerical control to a neural delta controller, the proposed architecture addresses key limitations of monolithic LLM-based control, including unstable trajectories, slow convergence, and hallucinated actions. The experimental results demonstrate that this decomposition leads to more stable and efficient closed-loop behavior.
Across all evaluated spatial tasks and locally deployed language models, the proposed LLM+DL framework consistently outperforms LLM-only baselines in convergence efficiency while preserving or improving task success rates. Average step reductions exceed 70%, with speedups of up to $8.83\times$, and these gains remain consistent even for smaller language models with weaker symbolic reasoning capabilities. These findings indicate that the observed performance improvements arise primarily from architectural design rather than language model scale, highlighting the effectiveness of neuro-symbolic integration for reliable and interpretable control.
Despite these advantages, the current study is limited to planar environments with perfect state observability and a fixed set of primitive spatial relations. The expressiveness of the symbolic layer is constrained by the predefined task vocabulary, and stochastic language model outputs and inference latency remain practical considerations. Future work will extend this framework to real-world robotic platforms, richer symbolic representations, hierarchical and compositional task structures, and vision-based environments, with the goal of enabling robust language-guided control in more complex and realistic settings.
@@ 6 Conclusion
Figure 10: Normalized distance-to-goal over time for all spatial tasks. The proposed LLM+DL framework converges faster and more stably than the LLM-only control across all relations.
@@
The efficacy of distinguishing symbolic reasoning from continuous execution in language-conditioned control tasks is further supported by these results.